# BOUQuET [Uncaptioned image]: dataset, Benchmark and Open initiative for Universal Quality Evaluation in Translation

### v1

@@ Abstract
This paper presents BOUQuET, a multicentric and multi-register/domain dataset and benchmark, and its broader collaborative extension initiative. This dataset is handcrafted in non-English languages first, each of these source languages being represented among the 23 languages commonly used by half of the world’s population and therefore having the potential to serve as pivot languages that will enable more accurate translations. The dataset is specially designed to avoid contamination and be multicentric, so as to enforce representation of multilingual language features. In addition, the dataset goes beyond the sentence level, as it is organized in paragraphs of various lengths. Compared with related machine translation (MT) datasets, we show that BOUQuET has a broader representation of domains while simplifying the translation task for non-experts. Therefore, BOUQuET is specially suitable for the open initiative and call for translation participation that we are launching to extend it to a multi-way parallel corpus to any written language.
@@ 1 Introduction
Although multilingual large-language model evaluation benchmarks are only starting (Dac Lai et al., 2023), there is a rich research history in multilingual evaluation datasets for natural language processing; e.g., Sun and Duh (2020); Malmasi et al. (2022); Yu et al. (2022), with Machine Translation (MT) being the task with the highest investment in multilinguality (Kocmi et al., 2024). This is evident from the nearly 15-year history of the Conference on Machine Translation (formerly a workshop, WMT), which has established an international evaluation campaign (Kocmi et al., 2024). The campaign has compiled a comprehensive collection of parallel data evaluations covering a broad range of language pairs, domains, and tasks. Although some datasets offer multi-way parallelism across languages (e.g., the TED parallel corpus), [^1] https://hal.science/hal-04702210v1 the largest multi-way parallel corpus was only recently introduced with FLORES-101 (Goyal et al., 2022), later expanded to FLORES-200 (NLLB-Team et al., 2022) and to 2M-FLORES (Costa-jussà et al., 2024).
However, even if great progress has been made, advances in multilingual applications for diverse languages are hindered by limited evaluations. Existing datasets and benchmarks fall short due to having an English-centric focus, a narrow register, compromised quality from automated construction and mining, limited language coverage, or a static nature, in addition to being prone to contamination.
Similarly, in parallel with the previous progress, there have been several initiatives that called for data annotation in a collaborative and open way, such as the translation data collection initiative reported in Singh et al. (2024).
This precisely sets the stage for the introduction of a multilingual evaluation dataset and a benchmark, BOUQuET, which additionally combines community efforts. The purpose of this paper is two-fold. First, the paper details how we develop the Source-BOUQuET dataset, which is the necessary stepping stone towards an open initiative. Second, it presents how we design the open initiative itself, which aims to build the Full-BOUQuET dataset; i.e., Source-BOUQuET translated into any written language.
1.1 Definitions
@@
We understand the term register as a functional variety of language that includes socio-semiotic properties, as expressed in Halliday and Matthiessen (2004), or more simply as a “contextual style”, as presented in Labov , pp.79–99). In that regard, a register is a specific variety of language used to best fit a specific communicative purpose in a specific situation.
@@ 2 Dataset: Source-BOUQuET
In this section, we describe the creation criteria that have follow to design Source-BOUQuET, as well as the languages it includes.
@@ 2.1 Main characteristics
As described in greater detail next, the Source-BOUQuET dataset is mainly characterized by its non-English-centric focus, its diverse registers and domain (which is complementary to FLORES-200), its manual and original composition, and its built-in dynamic extensibility.
@@
Source-BOUQuET is handcrafted by proficient speakers of French, German, Hindi, Indonesian, Mandarin Chinese, Russian, and Spanish. Each of those languages provides the same number of sentences to the final dataset. The languages for Source-BOUQuET (see Table 2 in Section 2.3) are all part of the top 20 languages in the world in terms user population, as listed in Eberhard et al. (2024). In addition, they are also used by a large number of both native and non-native speakers, which makes them good candidates for what we refer to as pivot languages; i.e., higher-resource languages that can facilitate—as source languages—the translation of datasets into lower-resource languages. English is often used as such a pivot language, since numerous people have a high degree of proficiency in English as a second language. English is not the only language in this situation, however, and is not always the best pivot language option. For example, it is much easier to find Guarani-Spanish bilingual speakers than it is to find Guarani-English bilingual speakers. What is more, cultural proximity may also make translation slightly easier.
@@
Registers derive from communicative purposes and, as such, are related to domains. However, the relationship between registers and domains is not one to one. For example, if we take a domain such as TV news, we can identify at least 3 registers:
  • The register used by the news anchor, which is represented by fully scripted language that is read from a teleprompter with a very specific and unnatural form of diction (hypercorrect enunciation, unnatural intonation, homogeneous pace).
  • The register produced by communication specialists (i.e., people who have been trained to be spokespersons or surrogates). The points they make have been scriptedand rehearsed to the point of being known by heart. Communication specialists also learn rehearsed patterns of language to make sure they do not get surprised by a question and can redirect the conversation towards the points they need to make instead of answering unexpected questions. All the language they use has been written and rehearsed. It sounds spontaneous but it is not structured like informal language.
  • The register represented in person-in-the-street segments, which is more informal and spontaneous (possibly colloquial).
The example above is taken from a domain where both speech and writing are used but the situation is not significantly different in the written modality only. Language users all commonly shift between registers, which is typically referred to as style-shifting. Style-shifting (i.e., register-shifting) occurs within domains; so the domain itself is not a fool-proof way of getting a specific register. Although the norms of the domain can impose the degree of formality and of lexical specialization, it is often the register (which derives from the communicative purpose), not the domain, that determines many aspects of linguistic structure (lexical density, pronoun use, syntax, etc.).
The registers we cover is Source-BOUQuET are given in Figure 1.
@@
To develop Source-BOUQuET, we set a variety of linguistic criteria that need to be covered. Guidelines are then shared with linguists who manually craft sentences covering examples of these linguistic criteria and compose paragraphs of various lengths.
The main strategies for open collaboration are to design contribution guidelines and build the annotation tool that enables to freely collect translations in any language. For this, BOUQuET is initially composed on 8 major languages plus the English translation and it is shared in a repository that allows language communities to easily add a new language even if they do not speak English, e.g. a native speaker of Quechua can translate from Spanish. This repository contains detailed guidelines on how to do it. BOUQuET’s innovative approach ensures universal language accessibility. This open collaborative initiative will enrich BOUQuET with the following characteristics.
@@ 2.2 Creation criteria
For the design of the creation guidelines, which are reported in detail in Appendix A, we prepared a list of linguistic coverage requirements as well as some statistical information.
@@
In order for BOUQuET to be representative of various linguistic phenomena, linguistic coverage requirements are defined (as listed in Table 1), which are to be included in sentences that form paragraphs. This continuity will include variation in the number of sentences per paragraph and in the lengths of sentences themselves.
@@
Different sentence structures (affirmation, interrogation, negation, subordination, coordination)
@@
Source-BOUQuET is intended to cover 8 domains: narration (as in fiction writing), dialog, social media posts, social media comments, how-to, miscellaneous website content (excluding social media or news), opinion pieces, miscellaneous other (such as written speeches or signage). The choice of these domains optimizes for variety and popular usefulness.
@@
Source-BOUQuET is built with register variety in mind, differently from FLORES-200, which covers a few different domains but remains largely within similar registers. We characterize the registers through 3 main aspects (connectedness, preparedness, and social differential), and a variety of options within these aspects (see Figure 1 for detail). Connectedness attempts to describe the type of interaction typically available in a given domain. Preparedness aims to gauge how much time is typically used to produce or edit language content. Social differential tries to describe various social situations in which language content is typically produced in each of the domains. Options are not mutually exclusive at the level of the domain but become differentiated at the level of a single sentence.
@@
Figure 1: Covered register characteristics.
Statistical guidance.
@@
In order to appropriately cover linguistic requirements and adequately represent the registers and domains, we performed a statistical analysis to understand each of the linguistic characteristics. In particular, our analysis covers most domains that we are including in Source-BOUQuET by using diverse public datasets: narration (Books3, gutenberg (Gerlach and Font-Clos, 2018)); Social media posts (Reddit (Baumgartner et al., 2020)); Social media comments (Wikipedia comments [^2] https://www.kaggle.com/competitions/jigsaw-multilingual-toxic-comment-classification); Conversations / Dialogues (dialogsum (Chen et al., 2021), Open Orca (Lian et al., 2023)); Tutorials/how-to articles (how-to Wikipedia-lingua  [^3] https://huggingface.co/datasets/GEM/wiki_lingua); Website content (c4 (Raffel et al., 2020)); News / Reflection pieces (cnndailymail (Nallapati et al., 2016), xsum (Narayan et al., 2018)) and Miscellaneous (Wikipedia). Note that we collect information from public data that do not always accurately match our categories but constitute a proxy. For each of these domains, we have analyzed dimensionality: characters per token; tokens per sentence and sentences per paragraph.
@@
Figure 2: (Left) Tokens per sentence and (middle) sentences per paragraph (Right) CEFR per domain.
Regarding tokens per sentence (Figure 2 left), we can see correlations between different domains, and clear differences in length, especially in dialogs which tend to be much shorter.
@@
Regarding sentences per paragraph (Figure 2 middle), we can find correlation between different datasets representing the same domain, where fiction writing paragraphs tend to be much longer (averaging 5 but reaching up to 20 ), dialogs and news articles are much shorter (barely reaching 3-4 sentences in a paragraph), and the rest of the categories are somewhere in between (normally staying between 1-5 but reaching up to 10 in some cases).
Furthermore, we have checked complexity: distribution of CEFR level [^4] https://rm.coe.int/common-european-framework-of-reference-for-languages-learning-teaching/16809ea0d4 as a proxy. Including here the % of C2 scores at sentence level for each dataset. This was labeled by a sonar-based model (Duquenne et al., 2023) trained on CEFR-SP data CEFR Text Classification which greatly outperformed Llama-3.1 (Touvron et al., 2023). See Figure 2 right.
Wikipedia seems to be the only dataset with a more considerable share of C2 sentences, with some others like dialogues having no samples scored as such.
Annotations.
Each entry of Source-BOUQuET includes the source text (in one of the 8 pivot languages from Table 2) and its translation into English, domain information, and contextual information for better translation accuracy.
@@ 2.3 Languages
BOUQuET aims to be as multicentric as possible, in contrast with most existing datasets like FLORES-200 that are English-centric. The motivation is mainly to be representative of linguistic phenomena. To this effect, it is created in 7 non-English languages. Each language contributes with a similar number of sentences along with their English equivalents given by the sentence creators themselves. See languages in Table 2.
ISO 639-3 ISO 15924 Language Family Subgroup1
cmn Hans Mandarin Chinese Sino-Tibetan Sinitic
deu Latn German Indo-European Germanic

### v2

@@ Abstract
BOUQuET is a multi-way, multicentric and multi-register/domain dataset and benchmark, and a broader collaborative initiative. This dataset is handcrafted in 8 non-English languages. Each of these source languages are representative of the most widely spoken ones and therefore they have the potential to serve as pivot languages that will enable more accurate translations. The dataset is multicentric to enforce representation of multilingual language features. In addition, the dataset goes beyond the sentence level, as it is organized in paragraphs of various lengths. Compared with related machine translation datasets, we show that BOUQuET has a broader representation of domains while simplifying the translation task for non-experts. Therefore, BOUQuET is specially suitable for crowd-source extension for which we are launching a call aiming at collecting a multi-way parallel corpus covering any written language. The dataset is freely available at https://huggingface.co/datasets/facebook/bouquet.
@@ 1 Introduction
Although multilingual large language model (LLM) evaluation benchmarks are only starting (Dac Lai et al., 2023), there is a rich research history in multilingual evaluation datasets for natural language processing; e.g., Sun and Duh (2020); Malmasi et al. (2022); Yu et al. (2022), with Machine Translation (MT) being the task with the highest investment in multilinguality (Kocmi et al., 2024). This is evident from the nearly 20-year history of the Conference on Machine Translation (formerly a workshop, WMT), which has established an international evaluation campaign (Kocmi et al., 2024). The campaign has compiled a comprehensive collection of parallel corpus evaluations covering a broad range of language pairs, domains, tasks and recently, investing in a multi-way parallel dataset expanding in languages Deutsch et al. (2025). However, the largest multi-way parallel evaluation dataset to date was introduced with FLORES-101 (Goyal et al., 2022), later expanded to FLORES-200 (NLLBTeam, 2024), FLORES+  [^1] https://oldi.org/ and to 2M-FLORES (Costa-jussà et al., 2024).
These existing datasets and benchmarks fall short due to having an English-centric focus, a narrow selection of registers, compromised quality from automated construction and mining, limited language coverage, or a static nature, in addition to being prone to contamination Sainz et al. (2023). Similarly, in parallel with the previous progress, there have been several initiatives that called for data annotation in a collaborative and open way, such as the translation data collection initiative Singh et al. (2024).
Recently, Wu et al. (2025) evaluate multilingual benchmarking and make a call for action for the need for accurate, contamination-free, challenging, practically relevant, linguistically diverse, and culturally authentic evaluations. This call and the urgent need of progressing in multilingual benchmarking set the stage for the introduction of a new multilingual multi-way parallel evaluation dataset and benchmark. BOUQuET, which additionally combines community efforts, relies on text written from scratch (contamination-free [^2] Note that BOUQuET is free from contamination in each initial state because it is originally created and not mined. However, from the moment we open-source certain splits, BOUQuET will risk to leak into training. Therefore, we keep one split hidden to avoid this.) by native speakers in 8 different major languages (linguistically diverse). Text includes a variety of 8 practical domains (practically relevant) that represent localised knowledge (culturally diverse). BOUQuET is aligned at the sentence and paragraph-level and it relies on a mixture of commissioned and openly collected human annotations to extend to any language.
The organisation of the paper is as follows. First, the paper details how we develop the Source-BOUQuET dataset (Section 3), which is the necessary stepping stone towards an open initiative. Second, we benchmark BOUQuET for the 8 pivot languages plus English (Section 4). Finally, Section 5 presents how we design the open initiative itself, which aims to build the Full-BOUQuET dataset; i.e., Source-BOUQuET translated into any written language. At the time of submission of this paper, BOUQuET includes 55 multi-way parallel completed languages (Table 6).
2 Definitions and background
Definitions
@@ 2 Definitions and background
We understand the term register as a functional variety of language that includes socio-semiotic properties, as expressed in Halliday and Matthiessen (2004), or more simply as a “contextual style,” as presented in Labov (1991, pp.79–99). In that regard, a register is a specific variety of language used to best fit a specific communicative purpose in a specific situation.
Background
There is a large body of work in creating datasets for MT evaluation (e.g. WMT International Evaluation Campaigns Deutsch et al. (2025)). However, the vast majority are limited to a few languages. we next discuss the main efforts to build massively multilingual MT benchmarks and one representation of multi-domain dataset.
FLORES+
FLORES+ (Maillard et al., 2024) is the largest multilingual extension of FLORES-200 Goyal et al. (2022) and it covers the largest multi-way parallel dataset in terms of languages in 3 domains (Wikipedia, News, Travel guides). Even if FLORES+ has paragraph information, the translation has been done at the level of sentence without showing context to the annotators.
NTREX-128
Similarly to FLORES+ NTREX-128 covers a multi-way parallel dataset but for 128 languages. Unlike FLORES-200, translators had the full context of the document available when translating sentences, but the authors did not know if (or to what extent) they used this information Federmann et al. (2022).
NLLB-MD
was motivated to complement FLORES-200 in terms of domains in the context of the NLLB NLLBTeam (2024) project. It covers chat, news and health domains in 6 languages and it includes a much larger number of sentences.
All these datasets are English-localised and English-centric, meaning that all languages have been translated from the original source English. They cover limited amount of domains (a maximum of 4) and do not differentiate among registers.
@@ 3 Dataset: Source-BOUQuET
In this section, we describe the creation criteria that have been followed to design Source-BOUQuET, as well as the languages it includes.
@@ 3.1 Main characteristics
As described in greater detail next, the Source-BOUQuET dataset is mainly characterized by its non-English-centric focus, its diverse range of registers and domains (which are complementary to FLORES-200), its manual and original composition, and its built-in dynamic extensibility. Table 1 provides a comparison of several relevant statistics from BOUQuET and the closest related datasets covered in the previous section.
@@
Source-BOUQuET is handcrafted by proficient speakers of Egyptian Arabic and MSA, French, German, Hindi, Indonesian, Mandarin Chinese, Russian, and Spanish. Each of these languages contributes the same number of sentences to the final dataset. The languages for Source-BOUQuET (see Table 3 in Section 3.3) are all part of the top 20 languages in the world in terms of user population, as listed in Eberhard et al. (2024). In addition, they are also used by a large number of non-native speakers, which makes them good candidates for what we refer to as pivot languages; i.e., higher-resource languages that can facilitate—as source languages—the translation of datasets into lower-resource languages. English is often used as such a pivot language, since numerous people have a high degree of proficiency in English as a second language. English is not the only language in this situation, however, and is not always the best pivot language option. For example, it is much easier to find Guarani-Spanish bilingual speakers than it is to find Guarani-English bilingual speakers. What is more, cultural proximity may also make translation slightly easier.
@@
Registers derive from communicative purposes and, as such, are related to domains. However, the relationship between registers and domains is not one to one. See the register and domain correspondence in Figure 5 (Appendix B). For example, if we take a domain such as TV news, we can identify at least 3 registers: (1) the register used by the news anchor, which is represented by fully scripted language that is read from a teleprompter with a very specific and unnatural form of diction (e.g., hypercorrect enunciation, unnatural intonation, homogeneous pace); (2) The register produced by communication specialists (i.e., people who have been trained to be spokespersons or surrogates). The points they make have been scripted and rehearsed to the point of being known by heart. It sounds spontaneous but it is not structured like informal language; (3) the register represented in person-in-the-street segments, which is more informal and spontaneous (possibly colloquial). This example is taken from a domain where both speech and writing are used but the situation is not significantly different in the written modality only. Language users all commonly shift between registers, which is typically referred to as style-shifting. Style-shifting (i.e., register-shifting) occurs within domains; so the domain itself is not a fool-proof way of getting a specific register. Although the norms of the domain can impose the degree of formality and of lexical specialization, it is often the register (which derives from the communicative purpose), not the domain, that determines many aspects of linguistic structure (e.g., lexical density, pronoun use, syntax, etc.).
@@
To develop Source-BOUQuET, we set a variety of linguistic criteria that need to be covered, including both unmarked and marked structures (e.g., expected and unexpected number agreement between subject and verb). Guidelines are then shared with linguists who manually craft sentences covering examples of these linguistic criteria and compose paragraphs ranging from 3 to 6 sentences in length. These paragraphs are then manually translated across all pivot languages.
The main strategies for open collaboration are to design contribution guidelines and build an annotation tool that enables the free collection of translations in any language. BOUQuET is shared in a repository that allows language community to easily add a new language by translating it from one of the 8 pivot languages or the English translation. This repository contains detailed guidelines on how to do it. BOUQuET’s innovative approach ensures widespread language accessibility. This open collaborative initiative will enrich BOUQuET with the following characteristics.
@@
Dataset Split ##\##Parag. ##\##Sent Avg. Wrd. Parag/Sent Reg. Dom. Lang. Dyn.
FLORES+ Dev $\times$× 997 Wikipedia, News, Travel guides 220
Devtest 1,012 25 $\times$×
Eval 992
NTREX-128 Test 123 1,997 389/24 $\times$× News 128 $\times$×
Dev 6,000
NLLB-MD Devtest $\times$× 1,310 25 $\times$× Chat, News, Health 6 $\times$×
Eval 1,500
BOUQuET Dev 120 504 Fiction, Conversation, Social media 55+ [^3] See Appendix D for language coverage details
Devtest 200 864 55/15 posts/comments, Tutorials, Website,
Eval 144 628 Reflection pieces, Miscellaneous
Table 1: Main statistics from MT evaluation datasets including BOUQuET: number of sentences, number of paragraphs, average word per paragraph (or sentence), register information, domains, languages, dynamism.
@@ 3.2 Creation criteria
For the design of the creation guidelines, detailed in Appendix A, we prepared a list of linguistic coverage requirements along with some statistical information.
@@
In order for BOUQuET to be representative of various linguistic phenomena, linguistic coverage requirements are defined (as listed in Table 2), which are to be included in sentences that form paragraphs. Sentences are assigned a unique identifier that combines a unique paragraph ID number with a serial sentence number. Thus, paragraphs can be retrieved by concatenating sentences that share the same paragraph ID.
@@
Different sentence structures (affirmation,
interrogation, negation, subordination, coordination)
@@
Source-BOUQuET is intended to cover 8 domains: narration (as in fiction writing), dialog, social media posts, social media comments, how-to manuals and instructions, miscellaneous website content (excluding social media or news), opinion pieces, and other miscellaneous (such as written speeches or signage). The choice of these domains optimizes for variety and popular usefulness.
@@
Source-BOUQuET is built with register variety in mind, differently from FLORES-200, which covers a few different domains but remains largely within similar registers. We characterize the registers through 3 main features (connectedness, preparedness, and social differential). Connectedness attempts to describe the type of interaction typically available in a given domain. Preparedness aims to gauge how much time is typically used to produce or edit language content. Social differential describes the relationship between the interlocutors involved in a given social situation (e.g., writer and reader, characters in a dialog, etc.). Each individual domain can present different combinations of features but become differentiated at the level of the sentence. There are a variety of feature combinations, which are mentioned in Figure 5 and defined in Appendix B.
@@
Statistical guidance for domain representation.
@@
In order to appropriately cover linguistic requirements and adequately represent domains, we performed a statistical analysis to understand the linguistic characteristics of each domain before creating BOUQuET. In particular, our analysis covers most domains that we are including in Source-BOUQuET by using diverse public datasets: narration (Books3, Gutenberg library (Gerlach and Font-Clos, 2018)); Social media posts (Reddit (Baumgartner et al., 2020)); Social media comments (Wikipedia comments [^4] https://www.kaggle.com/competitions/jigsaw-multilingual-toxic-comment-classification); Conversations / Dialogues (dialogsum (Chen et al., 2021), Open Orca (Lian et al., 2023)); Tutorials/how-to articles (how-to Wikipedia-lingua  [^5] https://huggingface.co/datasets/GEM/wiki_lingua); Website content (C4 (Raffel et al., 2020)); News / Reflection pieces (CNN-DailyMail (Nallapati et al., 2016), XSum (Narayan et al., 2018)) and Miscellaneous (Wikipedia). Note that we collect information from public data that do not always accurately match our categories but constitute a proxy. For each of these domains, we have analyzed dimensionality: characters per token; tokens per sentence and sentences per paragraph; and linguistic complexity with CEFR levels [^6] https://www.coe.int/en/web/common-european-framework-reference-languages/.
@@
Figure 1: (Top) Tokens per sentence and (Middle) sentences per paragraph (Bottom) CEFR per dataset representative of BOUQuET domains.
@@
Regarding tokens per sentence (Figure 1 left), we can see correlations between different domains, and clear differences in length, especially in dialogs which tend to be much shorter. Regarding sentences per paragraph (Figure 1 middle), we can find a correlation between different datasets representing the same domain, where fiction writing paragraphs tend to be much longer (averaging 5 but reaching up to 20 ), dialogs and news articles are much shorter (barely reaching 3-4 sentences in a paragraph), and the rest of the categories are somewhere in between (normally staying between 1-5 but reaching up to 10 in some cases).
To guide BOUQuET creators on the linguistic complexity required for each domain, we have assessed complexity using the distribution of CEFR levels as a proxy. This includes the % of C2 scores at the sentence level for each dataset. These scores were labeled by a SONAR-based model (Duquenne et al., 2023) trained on CEFR-SP data for CEFR Text Classification, which significantly outperformed Llama-3 (Touvron et al., 2023). See right side of Figure 1. Wikipedia seems to be the only dataset with a more considerable share of C2 sentences, with some others like dialogues having no samples scored as such.
Annotations and Quality Checks
Each entry of Source-BOUQuET includes the source text (in one of the 8 pivot languages of Table 3) and its translation into English, domain information and contextual information for better translation accuracy. To double-check that Source-BOUQuET does not contain repeated sentences, we explored the similarity across English sentences. For each English sentence, we computed SONAR embeddings Duquenne et al. (2023) and we computed the cosine distance on the vectors. There were only 14 sentences with a cosine distance below 0.3. These sentences are reported in Appendix E.
@@ 3.3 Languages
As mentioned earlier, BOUQuET aims to be multicentric and localized , in contrast to most existing datasets that are English-centric. The motivation is mainly to be representative of linguistic phenomena. To this effect, it is created in 8 non-English languages (Table 3). Each language contributes with a similar number of sentences along with their English equivalents given by the sentence creators themselves.
3.4 Multi-way extension to Source-BOUQuET languages
Details
Source-BOUQuET creators composed 250 sentences for each of the 8 pivot languages plus the corresponding English translation. The remaining 1,750 sentences for each pivot language are translated from English. The final Source-BOUQuET is composed of 2,000 sentences in 9 languages (8 pivot languages plus their translations into English).
Quality checks
Since multi-way parallel data is created from English, we manually checked that translations did not lose the linguistic information when translating from English. While translating BOUQuET, we had to make sure that the contextual information which was applicable to the whole paragraph was taken into consideration by the translators. To ensure this, we used a number of following QA strategies reported in Appendix C.
Additional contextual information
The multi-centric nature of BOUQuET is also a reminder that English is not morphologically rich (e.g., it doesn’t mark grammatical gender agreement between nouns, adjectives, and verbs) and displays relatively little information about formality in its written form (e.g., it uses only one second-person singular pronoun, regardless of who is addressing whom). As such, English isn’t an ideal source language for translation purposes unless translators can be provided with additional contextual information. The BOUQuET dataset includes such additional information; for example, the grammatical gender of the first and second person (when this isn’t obvious) or the linguistic markedness of some words or phrases (e.g., literary or archaic verb tenses, use of slang, infrequently used level of formality).
ISO ISO Language Family Subgroup1
6393 15924
arb Arab Modern Afro-Asiatic West Semitic
Standard Arabic
cmn Hans Mandarin Sino-Tibetan Sinitic
Chinese
deu Latn German Indo-European West Germanic